write an equation in slope intercept form: x-intercept 2, y-intercept -6
Answer by sachi(548)

an equation in slope intercept form: is y=mx+c
where m is the slope=6/2=3
& c is the y-intercept = -6
so the required eqn is y=3x-6
ans